# SA-Person: Text-Based Person Retrieval with Scene-aware Re-ranking

### v2

@@ A.4.1 SceneRanker Performance Across Retrieval Methods and Prompt Designs
     TBPS-CLIP [5]      50.29      75.24      81.58      61.54
     IRRA(ViT-B/16) [15]      60.57      83.35      87.86      70.65
     IRRA(ViT-L/14) [15]      63.37      84.54      88.67      72.71
     RDE [30]      61.26      83.50      88.23      71.17
      +NP      $77.13_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.87}}$      $86.83_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.33}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.87_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.70}}$
      +BOP      $76.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+15.03}}$      $86.88_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.38}}$      $88.23_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.46_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+10.29}}$
      +BEP      $\textbf{77.58}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+16.32}}$      $\textbf{86.84}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.34}}$      $\textbf{88.23}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $\textbf{82.18}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+11.01}}$
     RaSa [3]      62.51      83.83      88.09      72.05
      +NP      $77.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.53}}$      $86.66_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.83}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.76_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.71}}$
      +BOP      $76.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.73}}$      $86.74_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.91}}$      $88.09_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.29_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.24}}$
     MARS [10]      62.93      84.26      88.65      72.43
      +NP      $77.26_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+14.33}}$      $87.30_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+3.04}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $82.10_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.67}}$
      +BOP      $76.57_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+13.64}}$      $87.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.88}}$      $88.65_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$      $81.69_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+9.26}}$
To further assess SceneRanker’s generalizability, we evaluate its performance when integrated with various text-based person retrieval methods, each paired with three prompt designs: Naive Prompt (NP), Box Overlay Prompt (BOP), and Box Embedded Prompt (BEP). Results are presented in Table 6. SceneRanker significantly boosts performance for all methods, with R@1 improvements exceeding above 14.97% when using BEP, underscoring its compatibility and effectiveness in refining appearance-based retrieval. Across all methods, BEP consistently outperforms NP and BOP, achieving the highest R@1, suggesting that embedding bounding box coordinates enhances spatial grounding, enabling better alignment with textual descriptions.
A.4.2 Visualization of Retrieval Results
To illustrate the impact of SceneRanker, we present a qualitative comparison of retrieval results before and after applying SceneRanker, as shown in Figure 7.
@@ A.4.2 Visualization of Retrieval Results
Figure 7: Qualitative Comparison of Retrieval Results with and without SceneRanker. The leftmost column shows the query and ground-truth image. The green frame contains the retrieval results from the appearance-based stage before SceneRanker re-ranking, while the blue frame contains the retrieval results after SceneRanker re-ranking.
A.5 SceneRanker Prompt Design and Templates
This subsection details the instructional prompt templates used in the SceneRanker ablation studies. SceneRanker performs reranking without task-specific fine-tuning, using a multimodal prompt comprising a visual prompt $\mathbf{P}_{v}$ and an instructional textual prompt. Within each prompt strategy, $\mathbf{P}_{v}$ is fixed, but the instructional prompt varies to enhance spatial grounding, with multiple prompts tested. We describe: (1) the instructional prompt templates for three strategies evaluated for enhancing spatial grounding in SceneRanker, each in a dedicated table, and (2) the model-specific instructional prompt templates designed for Qwen-VL 2.5 [2], MiniCPM-V-2_6 [46], and InternVL 2.5 [8] in the MLLMs ablation study, tailored to each model’s capabilities while ensuring fair comparison. All prompts are designed for use with the InternVL2.5-8B model [8] unless otherwise specified.
A.5.1 Scene-aware Multimodal Prompt Templates
For each prompt strategy, the visual prompt $\mathbf{P}_{v}$ is fixed, but multiple instructional prompts were tested to optimize spatial grounding of pedestrian-specific region information. Below, we specify $\mathbf{P}_{v}$ for each strategy, describe the instructional prompt variations for reproducibility, and present examples in dedicated tables, with the final adopted prompt in bold.
[…]
"For the text description: ’text’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. Each image includes one person who is the intended candidate subject described in the text. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. In each image, one person is implicitly assumed to be the potential match for the individual described. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the given description: ‘$T$’, analyze the 10 provided candidate images. Each image contains one person implicitly corresponding to a potential subject referenced in the description. Evaluate how accurately each candidate aligns with the contextual details, visual cues, and semantic elements in the text description. Rank these images in descending order of relevance and contextual fidelity, with the most accurate match first. Return only a Python list of indices in the format: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7], where each number represents the original position of the image in the candidate set." "Given the description: ‘$T$’, 10 images are provided, each containing a person implicitly regarded as the candidate subject described in the text. Reorder the images based on how well they match the description, considering both the assumed subject’s appearance and the surrounding context. Provide the output as a Python list in descending order of relevance, formatted like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 7: Naive Prompt (NP) instructional prompt templates for SceneRanker.
"For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. Each image includes one person highlighted with a bounding box, representing the candidate subject mentioned in the text. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. In each image, a person is marked with a bounding box as the potential match for the individual described. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the given description: ‘$T$’, analyze the 10 provided candidate images. Each image contains one person highlighted with a bounding box, representing a potential subject referenced in the description. Evaluate how accurately each candidate aligns with the contextual details, visual cues, and semantic elements in the text description. Rank these images in descending order of relevance and contextual fidelity, with the most accurate match first. Return only a Python list of indices in the format: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7], where each number represents the original position of the image in the candidate set." "Given the description: ‘$T$’, 10 images are provided, each containing a person highlighted with a bounding box, representing the candidate subject described in the text. Reorder the images based on how well they match the description, considering both the highlighted subject’s appearance and the surrounding context. Provide the output as a Python list in descending order of relevance, formatted like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 8: Box Overlay Prompt (BOP) instructional prompt templates for SceneRanker.
"For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the text description: ‘$T$’, 10 images are provided. Each image includes one person marked by a bounding box (given as coordinates in <box>). This bounding box indicates the specific person selected as a potential match for the description. Please rank the images based on how accurately the boxed person and their surrounding scene match the description, starting with the most accurate match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the text description: ‘$T$’, 10 images are provided. Each image includes a bounding box, which indicates the specific person selected as a potential match for the description, and its coordinates are shown as <box> for each image. Please rank the images based on how accurately the boxed person and their surrounding scene match the description, starting with the most accurate match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. In each image, one person is implicitly assumed to be the potential match for the individual described. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." "For the description: ‘$T$’, 10 images are provided. Each image includes one person who is the intended candidate subject described in the text. Order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 9: Box Embedded Prompt (BEP) instructional prompt templates for SceneRanker.
[…]
A.5.2 Model-Specific Instructional Prompt Templates
For the MLLM ablation study, SceneRanker is evaluated with Qwen-VL 2.5, MiniCPM-V-2_6, and InternVL 2.5, each using a tailored instructional prompt template paired with plain full-scene images. These prompts are designed to leverage each model’s strengths while ensuring fair comparison. Their respective instructional prompt templates are provided in Table 10.
Qwen-VL 2.5: "For the text description: ‘$T$’, analyze the 10 candidate images and rank them from most to least consistent with the text. Output only a Python list containing all integers from 1 to 10, with no duplicates or omissions, like [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]." MiniCPM-V: "Considering the full context of the description ‘$T$’, rank the provided images from best to worst match. Output only a Python list containing numbers from 1 to 10." InternVL 2.5: "For the text description: ‘$T$’, order the images based on how accurately they reflect overall context in the text, starting with the most faithful match. Please output only a Python list format like this: [4, 2, 8, 1, 5, 3, 6, 9, 10, 7]."
Table 10: Instructional Prompt Templates for Comparative Evaluation of Multimodal Large Language Models.

### v3

@@ VI-E Versatility Across Diverse Retrieval Scenarios
IRRA (ViT-B/16) [jiang2023cross] 47.60 84.00 89.14 63.63
+SceneRanker $54.24_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+6.64}}$ $86.35_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.35}}$ $89.14_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $68.03_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.40}}$
IRRA (ViT-L/14) [jiang2023cross] 50.57 85.54 89.36 66.01
+SceneRanker $57.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+6.60}}$ $87.19_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+1.65}}$ $89.36_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $70.04_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+4.03}}$
RDE (ViT-B/16) [qin2024noisy] 48.73 83.78 88.88 63.93
+SceneRanker $56.51_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+7.78}}$ $6.17_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+2.39}} 88$$8.88_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+0.00}}$ $69.27_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\scriptsize+5.34}}$
TABLE VIII: SceneRanker Enhancements in General Image-Text Retrieval Scenarios.
Model (Backbone) Urban1k shareGPT4V
R@1 R@5 R@10 mAP R@1 R@5 R@10 mAP
CLIP (ViT-B/16) 53.10 78.40 85.20 64.34 47.60 73.92 83.12 59.72
+ SceneRanker 72.70 83.00 85.20 77.69 59.52 78.40 83.10 68.19
CLIP (ViT-L/14) 55.90 79.50 86.50 66.81 51.06 74.72 83.22 61.87
+ SceneRanker 73.10 83.90 86.50 78.36 61.34 78.94 83.22 69.53
@@
Figure 7: Qualitative Comparison of Retrieval Results with and without SceneRanker. The leftmost column shows the text and ground-truth image. The green frame contains the retrieval results from the appearance-based stage before SceneRanker re-ranking, while the blue frame contains the retrieval results after SceneRanker re-ranking.
Figure 8: Loglog Plot of Inference Time Across Gallery Sizes. The plot shows inference time (normalized units, log scale) versus gallery size NN (log scale) for three methods.
Table VII presents SceneRanker enhancements on augmented CUHK-SYSU in surveillance scenarios. It achieves R@1 gains ranging from 6.60% to 7.78% and mAP improvements from 4.03% to 5.34% on IRRA and RDE backbones, demonstrating the module’s robustness in complex surveillance environments. On SCENEPERSON-13W, SA-Person yields large improvements over baselines as shown in Tables IV. For Urban1K and ShareGPT4V, we integrate plug-and-play SceneRanker with CLIP baselines using full images and enriched descriptions. Table VIII reports retrieval performance on these datasets. SceneRanker enhances the baselines across both: R@1 improves by over 19% and mAP by more than 12% on Urban1K, effectively exploiting urban contextual cues such as landmarks and crowds. Comparable gains appear on ShareGPT4V, with $\sim$12% R@1 and $\sim$8% mAP uplifts, affirming adaptability to open-domain diversity. These results across scenarios confirm SA-Person’s broad applicability, extending beyond SCENEPERSON-13W to multimodal tasks requiring scene-aware disambiguation.
VI-F Computational cost of inference
Unlike prior approaches [niu2025chatreid] that rely on exhaustive MLLM-based retrieval across the entire gallery, SA-Person combines the efficiency of domain-specific models for appearance grounding with the robust contextual reasoning of MLLMs, achieving a strong balance between high retrieval performance and computational scalability. As defined in Section III, let $\mu_{s}$ and $\mu_{m}$ denote the per-image time units for domain-specific models and MLLMs, respectively, with $\mu_{m}\approx 2$ orders of magnitude larger than $\mu_{s}$ due to intensive multimodal processing. For a gallery of size $N$, Appearance-based Retrieval processes all $N$ full-scene images to yield $O(N\mu_{s})$ time overall, producing a compact set of $K$ candidates ($K=10\ll N$). Scene-aware Re-ranking then applies a single MLLM forward pass on these $K$ candidates with bounding boxes, incurring $O(K\mu_{m})$ time. The total time simplifies to $O(N\mu_{s}+K\mu_{m})$, orders of magnitude lower than gallery-wide MLLM retrieval at $O(N\mu_{m}+M\mu_{m})$ (where $K=10$, $M\ll N$), while surpassing pure domain-specific models at $O(N\mu_{s})$ through contextual re-ranking. As illustrated in Figure 8, this design renders the MLLM stage independent of $N$, ensuring scalability for large galleries.
VI-G Qualitative Results
To further validate the effectiveness of the proposed SA-Person framework, we present a qualitative comparison of retrieval results before and after applying the SceneRanker module in Figure 7. Without scene-aware re-ranking, the initial candidates retrieved by the appearance-based methods often include individuals with similar coarse attributes but in mismatched contexts, leading to suboptimal ranking. In contrast, SA-Person leverages an MLLM to interpret global scene semantics and re-rank these candidates accordingly. These qualitative examples demonstrate that SA-Person effectively bridges the gap between fine-grained appearance matching and holistic scene understanding, achieving more accurate and robust text-based person retrieval.
VII Limitations and Discussion
While SA-Person advances scene-aware text-based person retrieval, our approach has several limitations. First, the textual descriptions in SCENEPERSON-13W are generated via MLLMs to enable large-scale annotation, which efficiently captures diverse appearance and contextual details but may exhibit a relatively uniform stylistic phrasing compared to the diverse variations in real-world texts. A more effective strategy may involve incorporating diverse templates during generation to broaden stylistic coverage, and exploring directly optimizing the MLLM to select the most varied phrasing dynamically. Additionally, SceneRanker builds on off-the-shelf MLLMs in a training-free manner for plug-and-play flexibility and broad applicability, yet task-specific fine-tuning could amplify its precision in retrieval intent alignment and scene context disambiguation. Exploring lightweight adapters or domain-adaptive pretraining would offer a promising avenue to boost performance without sacrificing zero-shot robustness. These directions not only address potential refinements but also open avenues for extending SA-Person to interactive human-centric systems and dynamic video retrieval.
VIII Conclusion
In this paper, we tackle the challenging task of person retrieval by leveraging multimodal information from both pedestrian appearance and global scene context. Specifically, we make three key contributions. First, we present ScenePerson-13W, the first large-scale dataset offering comprehensive multimodal annotations that integrate fine-grained pedestrian appearance descriptions with rich scene-level contextual information within full-scene images. Second, we propose SA-Person, a two-stage retrieval framework that combines discriminative appearance grounding in cropped pedestrian regions with holistic scene understanding for robust retrieval from large-scale full-scene databases. Extensive experimental results demonstrate that SA-Person significantly outperforms existing baselines, underscoring the critical importance of integrating both local appearance cues and global scene context for robust and effective text-based person retrieval in complex real-world scenarios.
Acknowledgments
This work was supported in part by the InnoHK Program of the Hong Kong SAR Government; the National Natural Science Foundation of China under Grants No. 62306313, 62476188 and 62206276; the Natural Science Foundation of the Jiangsu Higher Education Institutions of China; and the Key Laboratory of New Generation Artificial Intelligence Technology & Its Interdisciplinary Applications (Southeast University), Ministry of Education, China.